# Achieving Near-Optimal Regret for Bandit Algorithms with Uniform Last-Iterate Guarantee

### v1

@@ Abstract
Existing performance measures for bandit algorithms such as regret, PAC bounds, or uniform-PAC (Dann et al., 2017), typically evaluate the cumulative performance, while allowing the play of an arbitrarily bad arm at any finite time $t$. Such a behavior can be highly detrimental in high-stakes applications. This paper introduces a stronger performance measure, the uniform last-iterate (ULI) guarantee, capturing both cumulative and instantaneous performance of bandit algorithms. Specifically, ULI characterizes the instantaneous performance since it ensures that the per-round regret of the played arm is bounded by a function, monotonically decreasing w.r.t. (large) round $t$, preventing revisits to bad arms when sufficient samples are available. We demonstrate that a near-optimal ULI guarantee directly implies near-optimal cumulative performance across aforementioned performance measures. To examine the achievability of ULI in the finite arm setting, we first provide two positive results that some elimination-based algorithms and high-probability adversarial algorithms with stronger analysis or additional designs, can attain near-optimal ULI guarantees. Then, we also provide a negative result, indicating that optimistic algorithms cannot achieve a near-optimal ULI guarantee. Finally, we propose an efficient algorithm for linear bandits with infinitely many arms, which achieves the ULI guarantee, given access to an optimization oracle.
@@ 1 Introduction
In online decision-making problems with bandit feedback, a learner sequentially interacts with an unknown environment: at each round, the learner plays an arm and then observes the corresponding reward of the played arm. Typically, the objective of the learner is to achieve a good cumulative performance, commonly measured by regret or probably approximately correct (PAC) bound. For instance, in the online advertisement scenario, the goal of the website (learner) could be maximizing the cumulative click numbers (Li et al., 2010). Hence, the website aims to minimize the regret that measures the cumulative clicks of the recommended ads compared to that of the unknown optimal ads. In addition to regret minimization, the goal could also be quickly identifying popular advertisements (Chen et al., 2014; Jin et al., 2019). To this end, a PAC bound is suitable here to measure the sample complexity (i.e., cumulative time steps) that the algorithm needs to identify those popular ads. To reap the benefits of both measures, Dann et al. (2017) propose a new performance measure, called uniform-PAC, ensuring that for all $\epsilon>0$, the total number of plays of $\epsilon$-suboptimal arms is bounded by a function, polynomial in $1/\epsilon$. The uniform-PAC bound can simultaneously imply a sublinear regret bound and a polynomial sample complexity for any desired accuracy.
Although the uniform-PAC provides a powerful framework to unify regret and PAC bound, it still fails to capture the instantaneous performance of the learning algorithm. In particular, a uniform-PAC algorithm could play a significantly bad arm for some large but finite round $t$, even if it enjoys a good cumulative performance. This drawback impedes the application of uniform-PAC algorithms into high-stakes fields. Clinical trials, for example, place high demands on instantaneous performance for every treatment test, since patients need to be assigned with increasingly better treatments when more experimental data are available (Villar et al., 2015). Hence, two natural questions arise:
@@
    Can we find a new performance measure that characterizes not only the cumulative performance but also the instantaneous performance?
@@
    If such a measure exists, is it optimally achievable by some bandit algorithms?
@@
In this paper, we answer both questions affirmatively. Specifically, our main contributions are summarized as follows:
@@
  • We introduce a new performance measure called uniform last-iterate (ULI), which simultaneously characterizes cumulative and instantaneous performance of bandit algorithms. On one hand, ULI can characterize the instantaneous performance: the per-round regret of any algorithm with ULI guarantee is upper bounded by a function, monotonically decreasing for large time $t$. On the other hand, we show that any algorithm with the ULI guarantee is also uniform-PAC, demonstrating that ULI can imply cumulative performance.
@@
  • To answer the question whether ULI is achievable, we examine three common types of bandit algorithms in the finite arm setting. First, we provide a stronger analysis to show that many existing elimination-based algorithms indeed enjoy the ULI guarantee. Second, we provide a lower bound for optimistic algorithms (such as UCB-type (Jamieson et al., 2014)), which shows that even if the optimistic algorithms enjoy a near-optimal cumulative performance, their instantaneous performance is significantly worse than elimination-based algorithms. Finally, we propose a meta-algorithm that enables any high-probability adversarial bandit algorithms, with a mild condition, to hold the ULI guarantee, and we demonstrate that many adversarial bandit algorithms naturally hold such a condition.
@@
  • If the number of arms is infinite, we propose an oracle-efficient [^1] A linear bandit algorithm is oracle-efficient if it calls an optimization oracle per-round at most polynomial times. linear bandit algorithm with the ULI guarantee (with access to an optimization oracle). In particular, we propose an adaptive barycentric spanner technique, selecting finitely many base arms that can represent all (possibly infinitely many) well-behaved arms. Leveraging the phased elimination algorithm (Lattimore et al., 2020), our algorithm can conduct the elimination over all arms by only playing a finite subset of arms and querying the optimization oracle only a polynomial number of times.
@@
Related Work. In bandit problems, regret and PAC bounds are widely adopted to evaluate the cumulative performance of algorithms. More concretely, one line of research (Auer et al., 2002a, b; Abbasi-Yadkori et al., 2011; Li et al., 2010; Zimmert and Seldin, 2019) aims to minimize the regret which measures the difference between the cumulative rewards of the selected arms and that of the best arm in hindsight. The PAC guarantees are more common than the regret when studying the pure-exploration/best arm identification problems (Even-Dar et al., 2006; Kalyanakrishnan et al., 2012; Zhou et al., 2014; Fiez et al., 2019; Li et al., 2022). One of the popular PAC measures is $(\delta,\epsilon)$-PAC which suggests that with probability at least $1-\delta$, the algorithm can output a near-optimal arm at most $\epsilon$ away from the optimal one by using a sample complexity polynomial in $1/\epsilon$. Later, Dann et al. (2017) introduce a new framework called uniform-PAC to unify both performance measures and develop an optimistic algorithm for episodic Markov decision processes (MDPs). Subsequent works design uniform-PAC algorithms for MDPs with linear function approximation (He et al., 2021) and bounded Eluder dimension (Wu et al., 2023). Though uniform-PAC strengthens regret and $(\delta,\epsilon)$-PAC bound, it still fails to characterize the instantaneous performance of bandit algorithms, in the sense that a uniform-PAC algorithm, even with a good cumulative performance, can play significantly bad arms for some large rounds.
@@
A seemingly related performance measure is last-iterate convergence (LIC) which has been studied for optimizing MDPs (Moskovitz et al., 2023; Ding et al., 2023) and they use the primal-dual approach to formulate the problem of identifying an optimal policy in the constrained MDPs from a game perspective. These works often require an exact knowledge of value functions and the LIC does not characterize the unknown dynamics of the environment. However, in our problem, the dynamics need to be learned as the algorithm sequentially interacts with the environment.
Notations. We use $e_{i}\in\mathbb{R}^{d}$ to denote a vector whose $i$-th coordinate is one and all others are zero. For any two vectors $x,y\in\mathbb{R}^{d}$, $x\leq y$ indicates that $x_{i}\leq y_{i}$ holds for each coordinate $i$. For a positive definite matrix $A\in\mathbb{R}^{d\times d}$, the weighted $2$-norm of vector $x\in\mathbb{R}^{d}$ is given by $\left\|{x}\right\|_{A}=\sqrt{x^{\top}Ax}$. For matrix $A=[a_{1},\cdots,a_{d}]\in\mathbb{R}^{d\times d}$ with each $a_{i}\in\mathbb{R}^{d}$, we use $A_{-i}$ to denote the $(d-1)$-tuple of vectors $[a_{1},\cdots,a_{i-1},a_{i+1},\cdot,a_{d}]$. For matrices $A,B$, we use $A\succ B$ to indicate that $A-B$ is positive definite. For two sets ${\mathcal{A}},{\mathcal{B}}$, we use ${\mathcal{A}}-{\mathcal{B}}$ to indicate the exclusion. For any scalar $C\in\mathbb{R}$, we use $\vec{C}$ to denote a vector, with all coordinates equal to $C$. For any $n\in\mathbb{N}$, let $[n]=\{1,\cdots,n\}$. We use $\widetilde{\mathcal{O}}(\cdot)$ to hide polylog factors.
@@ 2.1 Framework
We consider a general online decision-making framework where a learner interacts with a stochastic environment with a fixed arm set ${\mathcal{A}}$. At each round $t\in\mathbb{N}$, the learner plays an arm $A_{t}\in{\mathcal{A}}$ and observes the corresponding reward $X_{t,A_{t}}$. Assume that the rewards follow some stochastic patterns which will be specified for multi-armed bandits (MAB) and linear bandits, respectively as follows.
@@
Multi-armed bandits. In the stochastic MAB setting, the arm set follows that ${\mathcal{A}}=[K]$. Each arm $a\in[K]$ is associated with a fixed and unknown $[0,1]$-bounded distribution [^2] Note that all the MAB and linear stochastic bandit algorithms in this paper also work for $R$-subgaussian noise with minor adjustments. Only algorithms for adversarial bandits in Section 3.3 require the boundedness assumption. such that reward $X_{t,a}$ for each $t,a$ is an i.i.d. sample from this distribution with mean $\mu_{a}=\mathbb{E}[X_{t,a}]$. Let $\Delta_{a}=\mu^{\star}-\mu_{a}$ be the suboptimality gap where $\mu^{\star}=\max_{a\in[K]}\mu_{a}$.
@@
Linear bandits. In the stochastic linear bandits setup, we assume that the arm set ${\mathcal{A}}\subseteq\mathbb{R}^{d}$ is compact. The observed reward at round $t$ follows that $X_{t,A_{t}}=\left\langle{\theta,A_{t}}\right\rangle+\eta_{t}$ where $\theta\in\mathbb{R}^{d}$ is a fixed but unknown parameter, and $\eta_{t}$ is conditionally $1$-subgaussian. Let $\Delta_{a}=\sup_{b\in{\mathcal{A}}}\left\langle{\theta,b-a}\right\rangle$. We follow standard boundedness assumptions that $\left\|{\theta}\right\|_{2}\leq 1$, $\left\|{a}\right\|_{2}\leq 1$ for all $a\in{\mathcal{A}}$, and $\Delta_{a}\leq 1$ for all $a\in{\mathcal{A}}$.
@@
Suboptimality gap. For both MAB and linear bandits settings, we denote $a^{\star}$ as one of the optimal arms and use $\Delta=\inf_{a\in{\mathcal{A}}:\Delta_{a}>0}\Delta_{a}$ to denote the minimum suboptimality gap. Notice that it is possible that $\Delta=0$ when, for example, ${\mathcal{A}}$ is a ball. We use $\Delta_{A_{t}}$ to represent the instantaneous suboptimality, incurred by arm $A_{t}$ played at round $t$. For simplicity, we assume that the optimal arm is unique. Note that all our algorithms and analysis are applicable to the setting with multiple optimal arms.
2.2 Limitations of Existing Performance Measures
@@
Regret and $\left(\delta,\epsilon\right)$-PAC are widely adopted to measure the performance of a bandit algorithm. The regret is defined as:
@@
For any fixed $T\in\mathbb{N}$, the regret $R_{T}$ is defined as $R_{T}=\sum_{t=1}^{T}\Delta_{A_{t}}$.
Let $N_{\epsilon}=\sum_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{A_{t}}>\epsilon}\right\}}$ be the number of plays of arms whose suboptimality gap is greater than $\epsilon$. The definition of $\left(\delta,\epsilon\right)$-PAC is given as:
@@
[…]
As shown by Dann et al. (2017), both regret and $\left(\delta,\epsilon\right)$-PAC have limitations. Specifically, an algorithm with sublinear regret bound may play suboptimal arms infinitely often. For the algorithm with $\left(\delta,\epsilon\right)$-PAC guarantee, it may not converge to the optimal arm when feeding the algorithm with more samples. As a consequence, such an algorithm would play those arms with suboptimality gap, e.g., $\epsilon/2$ infinitely often. Motivated by these limitations, Dann et al. (2017) introduce uniform-PAC, defined as follows.
@@
[…]
We also call $F_{\text{UPAC}}(\delta,\epsilon)$ the sample complexity function. Uniform-PAC is a stronger performance measure than regret and $(\delta,\epsilon)$-PAC since it leads to the following implications, translated from MDPs to our settings.
@@
[…]
  • $\mathbb{P}\left(\lim_{t\to+\infty}\Delta_{A_{t}}=0\right)\geq 1-\delta$;
[…]
@@
Limitations of Uniform-PAC. According to Theorem 2.4, uniform-PAC can imply a long-term convergence (the first bullet) and good cumulative performance (the second and the third bullets), but it does not capture the convergence rate of $\Delta_{A_{t}}$ for each round $t$. In other words, even if an algorithm enjoys uniform-PAC, it could still play a significantly bad arm for some very large but finite $t$. This would lead to catastrophic consequences in safety-critical applications.

### v2

# Uniform Last-Iterate Guarantee for Bandits and Reinforcement Learning

@@ Abstract
Existing metrics for reinforcement learning (RL) such as regret, PAC bounds, or uniform-PAC (Dann et al., 2017), typically evaluate the cumulative performance, while allowing the agent to play an arbitrarily bad policy at any finite time $t$. Such a behavior can be highly detrimental in high-stakes applications. This paper introduces a stronger metric, uniform last-iterate (ULI) guarantee, capturing both cumulative and instantaneous performance of RL algorithms. Specifically, ULI characterizes the instantaneous performance by ensuring that the per-round suboptimality of the played policy is bounded by a function, monotonically decreasing w.r.t. round $t$, preventing revisiting bad policies when sufficient samples are available. We demonstrate that a near-optimal ULI guarantee directly implies near-optimal cumulative performance across aforementioned metrics, but not the other way around. To examine the achievability of ULI, we first provide two positive results for bandit problems with finite arms, showing that elimination-based algorithms and high-probability adversarial algorithms with stronger analysis or additional designs, can attain near-optimal ULI guarantees. We also provide a negative result, indicating that optimistic algorithms cannot achieve near-optimal ULI guarantee. Furthermore, we propose an efficient algorithm for linear bandits with infinitely many arms, which achieves the ULI guarantee, given access to an optimization oracle. Finally, we propose an algorithm that achieves near-optimal ULI guarantee for the online reinforcement learning setting.
@@ 1 Introduction
In online decision-making problems with bandit feedback, a learner sequentially interacts with an unknown environment: in each round, the learner plays an policy and then observes the corresponding rewards of the played policy. Typically, the goal of the learner is to achieve good cumulative performance, commonly measured by regret or probably approximately correct (PAC) bound. For instance, in the online advertisement scenario, the goal of the website (learner) could be maximizing the cumulative click numbers (Li et al., 2010). Hence, the website aims to minimize the regret that measures the cumulative clicks of the recommended advertisement compared to that of the unknown optimal advertisement. In addition to regret minimization, the goal could also be quickly identifying popular advertisements (Chen et al., 2014; Jin et al., 2019). To this end, a PAC bound is suitable here to measure the sample complexity (i.e., cumulative time steps) that the algorithm needs to identify those popular advertisements. To reap the benefits of both measures, Dann et al. (2017) propose a new performance measure called uniform-PAC, ensuring that for all $\epsilon>0$, the total number of $\epsilon$-suboptimal policies played by the algorithm is bounded by a function polynomial in $1/\epsilon$. The uniform-PAC bound can simultaneously imply a high-probability sublinear regret bound and a polynomial sample complexity for any desired accuracy.
Although uniform-PAC provides a powerful framework to unify regret [^1] Throughout the paper, we focus on high-probability regret, and therefore, when we mention regret, it always refer to high-probability regret. We refer readers to Remark 2.8 for a discussion on expected regret. and PAC bound, it still fails to capture the instantaneous performance of the learning algorithm. In particular, a uniform-PAC algorithm could play a bad policy in some late but finite round $t$, even if it enjoys a good cumulative performance. This drawback impedes the application of uniform-PAC algorithms into high-stakes fields. Clinical trials, for example, place high demands on instantaneous performance for every treatment test, since patients need to be assigned with increasingly better treatments when more experimental data are available (Villar et al., 2015). Hence, two natural questions arise:
@@
    Can we find a new metric that characterizes not only the cumulative performance but also the instantaneous performance?
@@
    If such a metric exists, is it optimally achievable by some algorithm?
@@
In this paper, we answer both questions affirmatively. Our main contributions are summarized as follows.
@@
  • We introduce a new metric called uniform last-iterate (ULI), which simultaneously characterizes cumulative and instantaneous performance of sequential decision-making algorithms. On one hand, ULI can characterize the instantaneous performance: the per-round suboptimality of any algorithm with ULI guarantee is upper bounded by a function, monotonically decreasing for late time $t$. On the other hand, we show that any algorithm with a (near-optimal) ULI guarantee is also (near-optimally) uniform-PAC, demonstrating that ULI can imply cumulative performance.
@@
  • To answer the question whether ULI is achievable, we examine three common types of bandit algorithms in the finite arm setting. First, we provide a stronger analysis to show that many existing elimination-based algorithms indeed enjoy a near-optimal ULI guarantee. Then, we propose a meta-algorithm that enables any high-probability adversarial bandit algorithms, with a mild condition, to achieve a near-optimal ULI guarantee, and we show that such condition naturally holds for many adversarial bandit algorithms. Finally, we provide a hardness result showing that optimistic algorithms (e.g., lil’UCB (Jamieson et al., 2014)) cannot achieve near-optimal ULI guarantee. As lil’UCB is near-optimally uniform-PAC, our hardness result also implies that ULI is strictly stronger than uniform-PAC.
@@
  • For linear bandits with infinitely-many arms, we propose an oracle-efficient [^2] A linear bandit algorithm is oracle-efficient if it calls an optimization oracle per-round for at most polynomial number of times. linear bandit algorithm with the ULI guarantee (with access to an optimization oracle). In particular, we propose an adaptive barycentric spanner technique, selecting finitely many base arms that can linearly represent all (possibly infinitely many) well-behaved arms. This technique generalizes the one in (Awerbuch and Kleinberg, 2008) for elimination-based algorithms by adaptively identifying spaces that active arms span. Leveraging the phased elimination algorithm (Lattimore et al., 2020), our algorithm can conduct the elimination over all arms by only playing a finite subset of arms and querying a linearly-constrained optimization oracle for only a polynomial number of times.
  • Finally, we propose a new algorithm for tabular episodic Markov decision processes (MDPs), which achieves a near-optimal ULI guarantee. In particular, our algorithm adapts uncertainty-driven reward functions to encourage exploration of the transition model, which ensures accurate estimations of value functions across all policies. The final ULI guarantee is achieved by conducting policy elimination.
@@
Related work. In online decision-making problems, regret and PAC bounds are widely adopted to evaluate the cumulative performance of algorithms. More concretely, one line of research (Auer et al., 2002a, b; Abbasi-Yadkori et al., 2011; Li et al., 2010; Jin et al., 2018) aims to minimize the regret which measures the difference between the cumulative rewards of the selected policies and that of the best policy in hindsight. The PAC guarantees are more common than the regret when studying the pure-exploration/best policy identification problems (Even-Dar et al., 2006; Kalyanakrishnan et al., 2012; Wagenmaker et al., 2022). One of the popular PAC measures is $(\delta,\epsilon)$-PAC which suggests that with probability at least $1-\delta$, the algorithm can output a near-optimal policy at most $\epsilon$ away from the optimal one by using a sample complexity polynomial in $1/\epsilon$. Later, Dann et al. (2017) introduce a new framework called uniform-PAC to unify both metrics and develop a uniform-PAC algorithm for episodic Markov decision processes (MDPs). Subsequent works design uniform-PAC algorithms for MDPs with linear function approximation (He et al., 2021) and bounded Eluder dimension (Wu et al., 2023). Though uniform-PAC strengthens regret and $(\delta,\epsilon)$-PAC bound, it still fails to characterize the instantaneous performance of online algorithms, i.e., a uniform-PAC algorithm, even with a good cumulative performance, can play bad policies for some late rounds.
@@
A seemingly related performance measure is last-iterate convergence (LIC) which has been studied for optimizing MDPs (Moskovitz et al., 2023; Ding et al., 2023) and they use the primal-dual approach to formulate the problem of identifying an optimal policy in the constrained MDPs from a game-theoretic perspective. These works often require additional knowledge of the value functions and the LIC does not characterize the unknown dynamics of the environment. However, in our problem, the dynamics need to be learned as the algorithm sequentially interacts with the environment.
@@ 2.1 Framework
We consider a general online sequential decision-making framework where a learner interacts with an environment with a fixed decision set. At each round $t\in\mathbb{N}$, the learner makes a decision from the set and observes the corresponding reward(s). In what follows, we instantiate this framework to multi-armed bandits, linear bandits, and tabular episodic Markov decision processes (MDPs).
@@
Multi-armed bandits. In the stochastic MAB setting, the arm (decision) set follows that ${\mathcal{A}}=[K]\triangleq\{1,\ldots,K\}$. Each arm $a\in[K]$ is associated with a fixed and unknown $[0,1]$-bounded distribution [^3] Note that all MAB and linear stochastic bandit algorithms in this paper also work for $R$-subgaussian noise with minor adjustments. Only adversarial bandits algorithms in Section 3.1 require the boundedness assumption. such that $\forall t$, reward $X_{t,a}$ is an i.i.d. sample from this distribution with mean $\mu_{a}=\mathbb{E}[X_{t,a}]$. Let $A_{t}$ be the arm played at round $t$ and $\Delta_{a}=\mu^{\star}-\mu_{a}$ be the suboptimality gap where $\mu^{\star}=\max_{a\in[K]}\mu_{a}$.
@@
Linear bandits. In the stochastic linear bandits setup, we assume that the arm set ${\mathcal{A}}\subseteq\mathbb{R}^{d}$ is compact. The reward of played arm $A_{t}$ at round $t$ follows that $X_{t,A_{t}}=\left\langle{\theta,A_{t}}\right\rangle+\eta_{t}$ where $\theta\in\mathbb{R}^{d}$ is a fixed but unknown parameter, and $\eta_{t}$ is conditionally $1$-subgaussian. Let $\Delta_{a}=\sup_{b\in{\mathcal{A}}}\left\langle{\theta,b-a}\right\rangle$. We follow standard assumptions that $\left\|{\theta}\right\|_{2}\leq 1$, $\left\|{a}\right\|_{2}\leq 1$ for all $a\in{\mathcal{A}}$, and $\Delta_{a}\leq 1$ for all $a\in{\mathcal{A}}$.
Tabular episodic MDPs. A tabular episodic MDP is formalized as ${\mathcal{M}}=({\mathcal{S}},{\mathcal{A}},H,r,P,\mu)$ where ${\mathcal{S}},{\mathcal{A}}$ are finite state and action spaces with $|{\mathcal{S}}|=S,|{\mathcal{A}}|=A$, $H$ is the horizon length, $r=\{r_{h}\}_{h=1}^{H}$ where $r_{h}:{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ is a known reward function, $\{P_{h}\}_{h\in[H]}$ where $P_{h}:{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$ is a transition function, and $\mu$ is the initial state distribution. At the beginning of each episode $t$, the learner executes a policy $\pi_{t}=\{\pi_{t,h}:{\mathcal{S}}\to\Delta({\mathcal{A}})\}_{h=1}^{H}$. Then, starting from the initial state $s_{t,1}\sim\mu$, for each stage $h\in[H]$, the learner repeatedly takes an action $a_{t,h}\sim\pi_{t,h}(s_{t,h})$, observes reward $r_{h}(s_{t,h},a_{t,h})$, and transits to the next state $s_{t+1,h}\sim P_{h}(\cdot|s_{t,h},a_{t,h})$.
For any policy $\pi$ and stage $h$, we define action value function $Q^{\pi}_{h}(s,a)$ and value function $V^{\pi}_{h}(s)$ as
$Q^{\pi}_{h}(s,a)=\mathbb{E}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,a_{h}=a,\pi\right],\ V^{\pi}_{h}(s)=\mathbb{E}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})\mid s_{h}=s,\pi\right].$
The optimal action value function and value function at each stage $h$ are denoted by $V_{h}^{\star}(s)=\max_{\pi}V_{h}^{\pi}(s)$, and $Q_{h}^{\star}(s,a)=\max_{\pi}Q_{h}^{\pi}(s,a)$ respectively. Let $\Delta_{\pi}=\mathbb{E}_{s_{1}\sim\mu}[V^{\star}_{1}(s_{1})-V^{\pi}_{1}(s_{1})]$.
@@
Suboptimality notations. For MAB and linear bandits settings, the instantaneous suboptimality is $\Delta_{t}=\Delta_{A_{t}}$, and for episodic MDP, $\Delta_{t}=\Delta_{\pi_{t}}$. We use $\Delta=\inf_{a\in\Pi:\Delta_{a}>0}\Delta_{a}$ to denote the minimum suboptimality gap. Notice that it is possible that $\Delta=0$ when, for example, arm set ${\mathcal{A}}$ is a ball.
2.2 Limitations of Existing Metrics
@@
Regret and $\left(\delta,\epsilon\right)$-PAC are widely adopted to measure the performance. The regret is defined as:
@@
For each fixed $T\in\mathbb{N}$, the regret $R_{T}$ is defined as $R_{T}=\sum_{t=1}^{T}\Delta_{t}$.
Let $N_{\epsilon}=\sum_{t=1}^{\infty}\mathbb{I}{\left\{{\Delta_{t}>\epsilon}\right\}}$ be the number of plays of policies whose suboptimality gap is greater than $\epsilon$. The definition of $\left(\delta,\epsilon\right)$-PAC is given as:
@@
For any fixed $\delta,\epsilon\in(0,1)$, an algorithm is $\left(\delta,\epsilon\right)$-PAC (w.r.t. function $F_{\text{PAC}}$) if there exists a function $F_{\text{PAC}}\left(\delta,\epsilon\right)$ polynomial in $\log(\delta^{-1})$ and $\epsilon^{-1}$ such that
$\mathbb{P}\left(N_{\epsilon}\leq F_{\text{PAC}}\left(\delta,\epsilon\right)\right)\geq 1-\delta.$
As shown by Dann et al. (2017), both regret and $\left(\delta,\epsilon\right)$-PAC have limitations. Specifically, an algorithm with sublinear regret bound may play suboptimal policies infinitely often. For the algorithm with $\left(\delta,\epsilon\right)$-PAC guarantee, it may not converge to the optimal policy when feeding the algorithm with more samples. Therefore, such an algorithm would play those policies with suboptimality gap, e.g., $\epsilon/2$ infinitely often. Motivated by these limitations, Dann et al. (2017) introduce uniform-PAC as follows.
@@
An algorithm is uniform-PAC for some fixed $\delta\in(0,1)$ if there exists a function $F_{\text{UPAC}}\left(\delta,\epsilon\right)$ polynomial in $\log(1/\delta)$ and $\epsilon^{-1}$, such that
$\mathbb{P}\left(\forall\epsilon>0:N_{\epsilon}\leq F_{\text{UPAC}}\left(\delta,\epsilon\right)\right)\geq 1-\delta.$
We also call $F_{\text{UPAC}}(\delta,\epsilon)$ the sample complexity function. Uniform-PAC is a stronger metric than regret and $(\delta,\epsilon)$-PAC since it leads to the following implications.
@@
If an algorithm is uniform-PAC for some $\delta$ with function $F_{\text{UPAC}}(\delta,\epsilon)=\widetilde{\mathcal{O}}\left(\alpha_{1}/\epsilon+\alpha_{2}/\epsilon^{2}\right)$ [^4] We use $\widetilde{\mathcal{O}}(\cdot)$ to hide polylog factors., where $\alpha_{1},\alpha_{2}>0$ are constant in $\epsilon$ and depend on $\log(1/\delta)$ and $K$ for MAB, $d$ for linear bandits, and $S,A,H$ for MDPs then, the algorithm guarantees:
  • $\mathbb{P}\left(\lim_{t\to+\infty}\Delta_{t}=0\right)\geq 1-\delta$;
  • $\left(\delta,\epsilon\right)$-PAC with $F_{\text{PAC}}(\delta,\epsilon)=F_{\text{UPAC}}(\delta,\epsilon)$ for all $\epsilon>0$;
@@
Limitations of Uniform-PAC. According to Theorem 2.4, uniform-PAC can imply a long-term convergence (the first bullet) and good cumulative performance (the second and the third bullets), but it does not capture the convergence rate of $\Delta_{t}$ for each round $t$. In other words, even if an algorithm enjoys uniform-PAC, it could still play a significantly bad policy for some very large but finite $t$. This would lead to catastrophic consequences in safety-critical applications.